If the activation energy for oxidation is 100 KJ mole^-1 , the ratio of oxidation rates at
800 degree Celcius and 500 degree Celcius is ?
1
Expert's answer
2020-06-16T09:36:33-0400

Solution.

"E_a=100KJ\/mol;"

"T_1=1073K;"

"T_2=773K;"

"k=e^{-\dfrac{E_a}{Rt}};"

"\dfrac{k_1}{k_2}=e^{-\dfrac{E_a}{R}(\dfrac{1}{T_1}-\dfrac{1}{T_2});}" ;

"\dfrac{k_1}{k_2}=e^{4.35}=75.24;"

Answer: "\dfrac{k_1}{k_2}=75.24."